# The Geometry of Queries: Query-Based Innovations in Retrieval-Augmented Generation

### v1

@@ Abstract
Digital health chatbots powered by Large Language Models (LLMs) have the potential to significantly improve personal health management for chronic conditions by providing accessible and on-demand health coaching and question-answering. However, these chatbots risk providing unverified and inaccurate information because LLMs generate responses based on patterns learned from diverse internet data. Retrieval Augmented Generation (RAG) can help mitigate hallucinations and inaccuracies in LLM responses by grounding it on reliable content. However, efficiently and accurately retrieving most relevant set of content for real-time user questions remains a challenge.
In this work, we introduce Query-Based Retrieval Augmented Generation (QB-RAG), a novel approach that pre-computes a database of potential queries from a content base using LLMs. For an incoming patient question, QB-RAG efficiently matches it against this pre-generated query database using vector search, improving alignment between user questions and the content. We establish a theoretical foundation for QB-RAG and provide a comparative analysis of existing retrieval enhancement techniques for RAG systems. Finally, our empirical evaluation demonstrates that QB-RAG significantly improves the accuracy of healthcare question answering, paving the way for robust and trustworthy LLM applications in digital health.
Keywords: RAG, LLM, Healthcare, Digital Health.
11footnotetext: $\dagger$ Equal contribution.
@@ 1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT (OpenAI, 2024), Meta’s Llama (Touvron et al., 2023) and Google’s Palm (Anil et al., 2023) and Gemini (Team et al., 2023) suite, have demonstrated remarkable capabilities across a diverse set of natural language understanding and generation tasks. This has led to several novel applications of language models across different domains. In healthcare, LLMs hold immense promise for developing conversational AI systems that can answer patient questions, offer personalized health advice, and potentially improve access to care, particularly for underserved populations (Clusmann et al., 2023; Peng et al., 2023; Thirunavukarasu et al., 2023; Alowais et al., 2023; Nori et al., 2023; Singhal et al., 2022, 2023; Tu et al., 2024).
@@
However, applying LLMs in healthcare presents significant challenges in ensuring the accuracy, reliability, safety, and adherence to the latest healthcare practices in the information they provide. The probabilistic nature of LLMs, combined with limitations and potential biases in their training data, can lead to the generation of fabricated or nonsensical information, known as hallucinations. This poses significant risks in a healthcare context where accurate information is paramount. Additionally, the knowledge embedded within an LLM is limited to the data it was trained on and may not reflect the most current medical guidelines and best practices. To address these challenges, further fine-tuning, instruction tuning (Wei et al., 2022) and additional reinforcement learning (Ouyang et al., 2022) approaches can be considered. Common approaches involve training the models on specific datasets containing the most up-to-date information and/or preferred responses. However, fine-tuning has its limitations: datasets can be difficult and expensive to acquire, especially in healthcare and the computation cost may be prohibitive.
@@
Retrieval Augmented Generation (RAG) (Lewis et al., 2020) offers a promising alternative by grounding LLM responses in a curated and reliable knowledge base of vetted information. Unlike relying solely on the LLM’s internal knowledge, RAG systems retrieve relevant information for a specific user query from this external knowledge base to inform the LLM’s answer generation process. This grounding helps to mitigate hallucinations, promotes factual accuracy, and allows developers to incorporate the latest knowledge. Applications using RAG systems have been shown to be quite compelling compare to traditional LLM fine tuning (Gupta et al., 2024; Ovadia et al., 2024). However, the effectiveness of RAG hinges critically on its ability to accurately retrieve the most pertinent information from the knowledge base, a task complicated by the inherent semantic gap between natural language user queries and the way information is structured and stored within a knowledge base (Ma et al., 2023). This “retrieval challenge” is a significant bottleneck for building effective RAG systems.
Several approaches have attempted to overcome this retrieval challenge by improving the alignment between user queries and the knowledge base. Some methods leverage the generative capabilities of LLMs themselves. For instance, Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2022) propose generating a hypothetical document that would ideally answer the user’s query. This synthetic document, more semantically aligned with the knowledge base content, is then used as the retrieval key. Similarly, QA-RAG (Kim and Min, 2024) takes a two-pronged approach: first generating a candidate answer using a fine-tuned LLM, and then leveraging both the initial query and this candidate answer to retrieve relevant information from the knowledge base. (Ma et al., 2023) proposes duplicating and/or splitting user queries and then retrieving content similar to either rewritten query. Alternatively, some methods focus on fine-tuning embedding models to create semantically aligned representations of queries and documents (Karpukhin et al., 2020). Despite these efforts, efficiently bridging the semantic gap between user queries and knowledge bases for accurate and reliable information retrieval in RAG systems remains an active area of research.
1.1 Contributions
This paper introduces a novel framework for analyzing and enhancing the retrieval phase of RAG systems. We propose that effective retrieval hinges on aligning user queries with the knowledge base across distinct semantic representations: query, answer, and content. To the best of our knowledge, we are the first to explicitly recognize these distinct “semantic spaces” for improving retrieval. While prior methods have explored alignment based on answer or content representations, our work proposes direct alignment within the query space itself. Existing approaches often implicitly combine these distinct representations, leading to sub-optimal retrieval. In contrast, we systematically disentangle these semantic spaces and demonstrate the advantages of our query based alignment strategy. Specifically, we make the following contributions:
@@ 1.1 Contributions
    QB-RAG: Query-Based Retrieval Augmented Generation: We propose QB-RAG, a novel retrieval method for RAG systems that addresses the misalignment between user queries and content bases by pre-generating a comprehensive set of questions from the knowledge base. Distinct from the methods that rely on generate hypothetical documents or candidate answers, QB-RAG directly maps incoming queries to this offline-generated question set using efficient vector search, enabling efficient, direct query-to-query comparison for enhanced retrieval.

### v2

@@ Abstract
Deploying Large Language Models (LLMs) for healthcare question answering requires robust methods to ensure accuracy and reliability. This work introduces Query-Based Retrieval Augmented Generation (QB-RAG), a framework for enhancing Retrieval-Augmented Generation (RAG) systems in healthcare question-answering by pre-aligning user queries with a database of curated, answerable questions derived from healthcare content. A key component of QB-RAG is an LLM-based filtering mechanism that ensures that only relevant and answerable questions are included in the database, enabling reliable reference query generation at scale. We provide theoretical motivation for QB-RAG, conduct a comparative analysis of existing retrieval enhancement techniques, and introduce a generalizable, comprehensive evaluation framework that assesses both the retrieval effectiveness and the quality of the generated response based on faithfulness, relevance, and adherence to the guideline. Our empirical evaluation on a healthcare data set demonstrates the superior performance of QB-RAG compared to existing retrieval methods, highlighting its practical value in building trustworthy digital health applications for health question-answering.
11footnotetext: † Equal contribution.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a diverse set of natural language understanding and generation tasks (OpenAI, 2024; Touvron et al., 2023; Anil et al., 2023). In healthcare, LLMs hold immense promise for developing conversational AI systems that can answer patient questions, offer personalized health advice, and improve access to care, particularly for underserved populations (Clusmann et al., 2023; Peng et al., 2023; Alowais et al., 2023; Nori et al., 2023; Singhal et al., 2023; Tu et al., 2024).
@@
However, applying LLMs in healthcare presents significant challenges in ensuring the accuracy, reliability, and adherence to the latest healthcare practices. The probabilistic nature of LLMs, combined with limitations and potential biases in their training data, can lead to hallucinations. Additionally, the knowledge within an LLM is limited to the data it was trained on and may not reflect latest clinical practice. To address these challenges, further fine-tuning, instruction tuning (Wei et al., 2022) and additional reinforcement learning (Ouyang et al., 2022) approaches can be considered. However, these post-training approaches still have limitations: datasets can be difficult and expensive to acquire, especially in healthcare, and the computation cost may be prohibitive.
@@
Retrieval Augmented Generation (RAG) (Lewis et al., 2020) offers a promising alternative by grounding LLM responses on a curated knowledge base of vetted information. Unlike relying solely on the LLM’s internal knowledge, RAG systems retrieve relevant information for a specific user query from this external knowledge base to inform the LLM’s answer generation. This grounding helps mitigate hallucinations and allows developers to incorporate the latest knowledge. Applications using RAG systems have been shown to be quite compelling compared to traditional LLM fine-tuning (Gupta et al., 2024; Ovadia et al., 2024). However, the effectiveness of RAG depends critically on its ability to accurately retrieve the most pertinent information from the knowledge base, a task complicated by the inherent semantic gap between user queries in natural language and the way information is structured and stored within a knowledge base (Ma et al., 2023). This “retrieval challenge” is a bottleneck for building effective RAG systems.
In this work, we tackle this problem with a new method called Query-Based Retrieval Augmented Generation (QB-RAG). Instead of matching a user’s query directly to documents, QB-RAG first transforms the knowledge base into a high-quality set of answerable questions. The user’s query is then matched to the most relevant pre-generated question to find the best context. Our core contribution is an automated, LLM-powered filter that vets each generated question for clarity and answerability, ensuring a reliable, query-centric knowledge base. This offline preparation is shown to significantly improve both retrieval and answer quality, providing a practical blueprint for more dependable RAG systems in healthcare.
Generalizable Insights about Machine Learning in the Context of Healthcare
Our work offers several generalizable insights for developing, evaluating, and deploying Large Language Model (LLM) applications in safety-critical domain of healthcare:
  • LLMs may Enable Automated, Quality-Controlled Curation of Knowledge Base for RAG: LLMs can be employed to automatically generate diverse questions from trusted healthcare content and may subsequently act as an ’answerability filter’ to ensure generated questions are relevant and verifiable against the source. This automated offline pipeline can create a high-quality, query-centric representation of knowledge base, enhancing grounding and reducing manual curation efforts often needed for reliable healthcare applications.
  • Comprehensive Evaluation Framework for Healthcare QA: We introduce a comprehensive evaluation framework specifically designed for healthcare question-answering. This framework goes beyond standard retrieval metrics by incorporating measures of answer quality, including faithfulness, relevancy, and adherence to healthcare guidelines. This allows for a more nuanced assessment of how retrieval improvements translate into tangible benefits for downstream question-answering in a healthcare context.
  • RAG Performance is Sensitive to Offline Question Set Quality: Our findings indicate that the effectiveness of query-based RAG approach, particularly its ability to generalize to the variability of real-world user health queries, is correlated with the coverage and diversity of the pre-generated question set. Practitioners can optimize offline question generation to capture a wide range of answerable intents from a knowledge base for effective real-world performance based on their unique content properties.
  • Offline Query-Space Alignment Combines Low Inference Latency with Enhanced Retrieval: Common RAG enhancement techniques involving multiple online LLM calls can introduce inference latency. Our work demonstrates that offline query-space alignment offers a practical solution. Pre-computing a query-centric knowledge representation can significantly reduce response times, making performant and reliable Healthcare-RAG systems more viable.
2 Related Work
To address this “retrieval challenge”, one set of approaches leverage the LLMs to create queries more semantically aligned with the knowledge base content. Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2022), for example, generate a hypothetical document that would ideally answer the user’s query. This synthetic document is then used as the retrieval key, improving the chances of finding relevant information. Similarly, QA-RAG (Kim and Min, 2024) utilizes a fine-tuned LLM to generate a candidate answer, which is then used in conjunction with the original query to enhance retrieval. Another set of approaches diversifies search space by modifying user queries, such as duplicating or splitting queries and then using these queries to retrieve wider range of relevant documents (Ma et al., 2023). Methods like Dense Passage Retrieval (Karpukhin et al., 2020) focus on embedding fine-tuning, aiming to create semantically aligned representations of queries and documents, hoping to enhance retrieval precision. Other work has focused on structuring the knowledge base itself, for instance by using knowledge graphs. Xu et al. (Xu et al., 2024) construct a graph from documents and retrieve relevant subgraphs to augment the generation process in customer service contexts.
Offline synthesis of pseudo-queries has also been explored for improving information retrieval. In doc2query and its iteration docTTTTTquery, (Nogueira et al., 2019b, a) showed that generating multiple queries from a document and concatenating them alongside the original document improves retrieval performance by enriching the document representation. More recently, (Raina and Gales, 2024) proposed a “Question-Based Retrieval using Atomic Units” method for enterprise RAG applications. They decompose knowledge into question-answer pairs based on atomic units of information, advocating for a question-centric retrieval approach. Concurrent to our work, QuIM-RAG (Saha et al., 2024) also generates a representative question for each document, performing retrieval by finding the closest match in an ”inverted question” index. Although these approaches share a similar high-level goal, they do not present scalable and robust query base filtering mechanisms, or they lack in comprehensive metrics at all levels of knowledge base curation, retrieval and answer generation, which are crucial for performance and error tracing in high stakes domains like healthcare. Our proposed method query-based RAG (QB-RAG) shares the spirit of these offline query generation approaches, but differentiates itself by focusing on generating answerable questions tailored for healthcare question-answering. This emphasis on answerability enforced by our LLM-based filtering mechanism is a crucial distinction, particularly in healthcare, where verifiable answers are paramount.
3 Health Application and Dataset
Digital health programs represent a paradigm shift in managing chronic conditions such as Type 2 diabetes (T2D) and Hypertension (HTN). They can improve health markers associated with chronic conditions by delivering personalized care and support directly to patients Majithia et al. (2020). An integral component of these digital health programs is the messaging platform. However, the increasing demand for immediate healthcare information often strains the capacity of human healthcare providers to deliver timely support through messaging.
To overcome this scalability challenge, digital health platforms are increasingly turning to meticulously curated content repositories designed to provide readily available answers to common patient questions. This study utilizes such a content repository, referred to as “Content Cards,” developed to address frequently asked questions (FAQs) from patients with T2D and HTN participating in our proprietary digital health programs. This dataset, encompassing 630 English-language content cards, covers a comprehensive range of topics pertinent to managing T2D and HTN. This includes general health information, detailed guidance on using platform-specific features and connected devices (e.g., blood glucose monitors), and personalized dietary recommendations. Figure 1 illustrates how these Content Cards are presented within our mobile application, while the excerpt in Appendix D provides a representative sample of the content.
Figure 1: Illustration of Content Cards in our mobile application, these cover broad range of topics, from operational health management to nutrition and physical advice.
4 Mathematical Formulation
This section presents a mathematical framework for analyzing retrieval within RAG systems, highlighting the limitations of conventional approaches and motivating our proposed QB-RAG method.
4.1 Notation
Consider the set of $M$ content documents $\mathcal{C}=\{c_{1},\cdots,c_{M}\}$ and the set of $N$ questions $\mathcal{Q}=\{q_{1},\cdots,q_{N}\}$ generated from $\mathcal{C}$ as specified in 5.1. Unless otherwise specified, $c\in\mathcal{C}$ and $q\in\mathcal{Q}$ will refer to their respective embedding representations. We will assume that all embeddings are normalized and the distance function is the cosine distance denoted as $d(x,y)$.
The question generation process in section 5.1 is a one to many process, that is from one content we generate multiple questions. To encode this relationship, we denote the matrix $A\in\{0,1\}^{M\times N}$ where $A_{ij}=\mathds{1}[c_{i}\text{ generated }q_{j}]$, where $\mathds{1}$ denotes the indicator function. Per our generation process and the relevance evaluation, it is understood that $A_{ij}=1\Rightarrow q_{j}\text{ can be answered using }c_{i}$. However, the converse is not necessarily true as potentially different contents may be able to answer the same question. Thus, we define $A^{*}\in\{0,1\}^{M\times N}$, the ground-truth answerability matrix, which is the dense unobserved matrix such that $A^{*}_{ij}=\mathds{1}[q_{j}\text{ can be answered using }c_{i}]$. This implies that $A$ is a partial observation of $A^{*}$, which we call the oracle matrix. Specifically, $A_{ij}=1$ if $A^{*}_{ij}=1$ and $q_{j}$ was generated from $c_{i}$ during the offline question generation process; otherwise, $A_{ij}=0$. We point out that following our question generation process, we have $\forall c_{j}\in\mathcal{C}:\>\exists q_{i}\in\mathcal{Q}\text{ s.t. }A_{ij}=1$.
For a new user question $q_{0}$, traditional RAG systems retrieve content that maximize some measure of similarity $\arg\max_{c\in\mathcal{C}}{c^{t}q_{0}}$ (or equivalently $\arg\min_{c\in\mathcal{C}}d(c,q_{0})$).
For simplicity in our mathematical formulation, we will assume retrieval to mean a single piece of content in this section, where RAG retrieves the most similar content to the user query. We use Google’s textembedding-gecko in the current work for generating embeddings. We denote the embedding matrices for content base and questions, respectively, with capital letters $C\in\mathbb{R}^{d\times M}$ and $Q\in\mathbb{R}^{d\times N}$ where $d$=768 is the dimension of our chosen embedder.
4.2 Ideal Retrieval Objective
Rather than pursuing this exact operation for every entry, many RAG systems typically approximate the retrieval task with some proxy as briefly described in 2. We briefly summarize them here in terms of our mathematical formulation.
  • Naive RAG: Approximates $f^{*}(c,q_{0})\approx c^{t}q_{0}$. Relies on cosine similarity between query and content embeddings, which can suffer from semantic misalignment. See a failing example in Appendix C.
  • LLM-Rewriting Methods (HyDE, Query2Doc, QA-RAG): Approximate
    $f^{*}(c,q_{0})\approx c^{t}p_{LLM}(q_{0})$, using LLMs to rewrite queries to better align with content space. Here, $p_{LLM}(q_{0})$ refers to the LLM’s output conditioned on the original query aligned with the content space. The effectiveness of this approach is highly dependent on the quality of LLM’s generation and alignment with the content base embeddings.
  • Adapter/Fine-tuning Methods: Approximate $f^{*}(c,q_{0})$ by learning transformations or specialized embeddings to improve alignment. These often require expensive training data and may not generalize well to different datasets.
4.3 QB-RAG
QB-RAG leverages a pre-computed set of questions, $\mathcal{Q}$, derived from the content base $\mathcal{C}$. These questions are generated offline as described in Section 5.1, mitigating concerns about online computational overhead. In contrast to methods relying on online LLM calls for query rewriting, QB-RAG shifts this computational burden offline. This distinction is significant, as online rewriting necessitates serial LLM invocation, potentially introducing latency detrimental to user experience. Furthermore, QB-RAG’s direct alignment within the query space offers a more transparent and interpretable retrieval process compared to the implicit alignment strategies of LLM-based rewriting techniques.
4.3.1 Vanilla QB-RAG
Algorithm 1 Vanilla QB-RAG
[…]
 Similarities $z\leftarrow Q^{t}q_{0}\in\mathbb{R}^{N}$
${}\mathcal{S}\leftarrow\{\}$
for $j\in[1..N]$ do
  $i\leftarrow i$ s.t. $A_{ij}=1$ identify the content associated associated to the $j$-th question.If $A$ is non-sparse, we break ties by favoring new content, by number of associated questions, then at random.
  $\mathcal{S}\leftarrow\mathcal{S}\cup\{c_{i}\}$ if $c_{i}\notin\mathcal{S}$
  If $\lvert\mathcal{S}\rvert=k$ then break
end for
Return $\mathcal{S}$
4.3.2 Oracle and Approximate QB-RAG
In this section, we present oracle and approximate variants as conceptual extensions. We leave experimental evaluation of these variants as future work.
We turn our attention to the best possible retrieval system using a query based alignment and retriever. We previously noted that the matrix $A$ indicating which content generated which question, is actually a sparse partial observation of $A^{*}$ indicating which content is relevant to a question. For this section we will assume that we have access to the oracle matrix $A^{*}$ or some approximation of it denoted $\hat{A}$. For simplicity, we describe the algorithms with ${A}^{*}$, but is also applicable to $\hat{A}$.